solve by elimination method.
7r-6s=-22
6r-7s=54
If you can provide the break down of this that would be greatly appriciated. I would like to use this in my notes for examples..Thanks A lot.

Answer by checkley77(12844)

7r-6s=-22 MULT. BY -6
6r-7s=54 MULT. BY 7 & ADD.
-42R+36S=132
42R-49S=378
----------------
-13S=510
S=510/-13
S=-510/13 ANS.
7R-6*-510/13=-22
7R+3060/13=-22
7R=-22-3060/13
7R=(-22*13-3060)/13
7R=(-286-3060)/13
7R=-3346/13
R=(-3346/13)*1/7
R=-3346/91 ANS.